# LAURA: Enhancing Code Review Generation with Context-Enriched Retrieval-Augmented LLM

### v1

@@ III-C Baselines
Tufano et al. [tufano2024code] compared ChatGPT-3 with T5CR [tufano2022using], CodeReviewer [li2022automating], and CommentFinder [hong2022commentfinder] in code review generation/recommendation, concluding that SOTA methods outperform ChatGPT-3. Among these methods, CodeReviewer is also designed for reviewing code diffs. Therefore, we choose CodeReviewer as a baseline to evaluate our method against traditional pre-trained models. Additionally, since LAURA is based on ChatGPT-4o and DeepSeek v3111We exclude DeepSeek R1 due to insufficient data after its release date., we also select ChatGPT-4o and DeepSeek v3 as additional baselines to evaluate the performance improvements introduced by our method.
@@ III-D Evaluation
Given the nature of LLMs and the diversity of natural language [fiske1992but], it is impractical to expect LLM outputs to closely match human reviewer comments. Thus, traditional metrics like BLEU and ROUGE are not suitable [reiter2018structured]. To efficiently evaluate the overall performance of the review comments generated by LAURA, we adopt an LLM-based evaluation approach. However, LLM evaluation has limitations when it comes to assessing the practical usefulness of review comments in real-world scenarios. Therefore, we also introduce a human evaluation approach.
@@ III-D1 LLM Evaluation
We follow the LLM-based DeepCRCEval method by Lu et al. [lu2025deepcrceval], with some modifications: (1) merging some metrics to improve clarity; (2) instructing the LLM to rate each comment on a 1–5 scale according to five metrics (higher scores indicating better quality), instead of the original 10-point scale, since prior study [joshi2015likert] shows that a 5-point scale is easier to understand and apply accurately, avoids the coarseness of a 3-point scale, and provides greater stability and consistency than a 10-point scale; (3) adding simple scoring guidelines for each metric. For example, for the “Readability” metric, we provided the LLM with the following explanation: “In the 1-5 scoring system, a score of 5 means the text is completely clear and readable; a score of 3 means the expressed meaning can be basically understood; and a score of 1 means the text is nearly unreadable or incomprehensible.” We believe these scoring guidelines help calibrate scores across different evaluation samples and positively impact the consistency of the LLM. Table II shows the final metrics.
@@ III-D2 Human Evaluation
Sometimes, LLM-based evaluation may not effectively assess the usefulness of the generated review comments in real-world scenarios. Therefore, we also conduct a human evaluation. Following Mahajan et al. [mahajan2020recommending], we adjust their method for LLM-generated results (Table III): we remove the “Unavailable” category, redefine the other three to suit our goals, and add an “Uncertain” category. This is because, in our evaluation, all test data have corresponding generated review comments, so the “Unavailable” case does not apply. However, given the limited knowledge of the repository and PR context, these review comments could be difficult to evaluate, leading to “Uncertain” cases. Specifically, since the data comes from 1,807 open-source projects, we cannot fully understand every project’s code requirements, organizational structure, or other contextual details. Therefore, even if we determine that the LLM-generated code review comments are generally useful, without factual grounding, we cannot be certain whether they would be accepted within the specific project context of the reviewed code. We categorize each generated review comment as “Instrumental”, “Helpful”, “Uncertain”, or “Misleading”, and use three metrics to evaluate quality.
@@ IV-A RQ1: Effectiveness of LAURA
In addition, in our experiments, the performance of directly using ChatGPT-4o for code review generation even surpasses that of CodeReviewer, which contrasts with the results of Tufano et al. [tufano2024code]. We attribute this to advances in LLMs since ChatGPT-3, dataset differences, and varied evaluation methods. Our dataset, possibly containing more long code diffs, may have overwhelmed pretrained models. Additionally, Tufano et al. evaluated the three traditional methods as a whole, which also leads to differences in results.
@@ VI-A Code Review Automation
Early studies mainly focused on reviewer recommendation, using methods such as file path analysis [thongtanunam2015should], traceability graphs [sulun2019suggesting], load balancing [asthana2019whodo], collaborative filtering [chueshev2020expanding], and balancing expertise and workload [mirsaeedi2020mitigating]. Other approaches considered cross-project collaboration [kong2022recommending] and composite analyses of file paths, project details, and author info [pandya2022corms]. Some research also aimed to predict which parts of the submitted code would need comments [hellendoorn2021towards].
Code review automation has advanced to include tasks like review necessity prediction, comment generation, and code refinement. Notable works include T5CR [tufano2022using], CodeReviewer [li2022automating], and CommentFinder [hong2022commentfinder]. T5CR, based on the Text-To-Text Transfer Transformer (T5) model [raffel2020exploring], handles method-level code for pre-review optimization, comment generation, and post-review refinement. CodeReviewer uses T5 architecture with CodeT5 initialization [wang2021codet5], processes line-level diffs, and supports review necessity prediction, comment generation, and code refinement across nine languages. CommentFinder is an information retrieval (IR)-based method that uses Bag of Words (BoW) and similarity measures to find similar code snippets and recommend relevant review comments.
So far, several works have attempted to apply LLMs to code review automation tasks. Lu et al. [lu2023llama] used the LLaMA model for review necessity prediction, comment generation, and code refinement, achieving results comparable to pre-trained models via parameter-efficient fine-tuning. Guo et al. [guo2024exploring] showed that ChatGPT-3.5 outperformed CodeReviewer in post-review refinement. Pornprasit et al. [pornprasit2024fine] explored ChatGPT-3.5’s performance in both pre- and post-review refinement, analyzing few-shot learning effects. Tufano et al. [tufano2024code] compared ChatGPT-3 with three existing methods and found it achieved state-of-the-art performance in code refinement, but lagged in review generation. However, these studies provide only a limited exploration of LLMs and have not fully unlocked their potential for code review automation tasks.
@@ VI-B Large Language Model
Language models have evolved from early statistical models to neural, pre-trained models, and now to large language models (LLMs) with over 10B parameters, trained on vast datasets [zhao2023survey]. LLMs show superior performance and emergent capabilities [wei2022emergent], often matching or surpassing task-specific pre-trained models. Notable LLMs include ChatGPT-4o, which excels across benchmarks [openAIGPT4o], and DeepSeek v3 [liu2024deepseek] among open-source models. Our research builds on ChatGPT-4o and DeepSeek v3, applying the LAURA framework to both.
So far, LLMs have been widely used in software engineering tasks like vulnerability prediction [sun2024gptscan, zhou2024large, sagodi2024reality] and commit message generation [eliseeva2023commit, wu2024commit, li2024only]. Many studies have explored ways to improve the performance of LLMs on specific tasks. Generally, LLMs can be further enhanced through methods such as augmenting input information [lu2024exploring], retrieval-augmented generation [gao2023retrieval], and prompt engineering [white2024chatgpt]. Augmenting input information is based on the intuitive idea that providing the model with richer contextual information helps it better understand the task and produce better results. Retrieval-augmented generation [lewis2020retrieval] enhances the model’s task awareness by incorporating knowledge from external databases, thereby strengthening the model’s generation capabilities. Prompt engineering [liu2023pre] involves designing prompts for the model, guiding it through instructions on the actions it must perform with the given input, and ensuring that the LLMs achieve high performance in downstream tasks.

### v2

@@ III-C Baselines
Tufano et al. [52] compared ChatGPT-3 with T5CR [53], CodeReviewer [27], and CommentFinder [20] in code review generation/recommendation, concluding that SOTA methods outperform ChatGPT-3. Among these methods, CodeReviewer is also designed for reviewing code diffs. Therefore, we choose CodeReviewer as a baseline to evaluate our method against traditional pre-trained models. Additionally, since LAURA is based on ChatGPT-4o and DeepSeek v3111We exclude DeepSeek R1 due to insufficient data after its release date., we also select ChatGPT-4o and DeepSeek v3 as additional baselines to evaluate the performance improvements introduced by our method.
@@ III-D Evaluation
Given the nature of LLMs and the diversity of natural language [13], it is impractical to expect LLM outputs to closely match human reviewer comments. Thus, traditional metrics like BLEU and ROUGE are not suitable [41]. To efficiently evaluate the overall performance of the review comments generated by LAURA, we adopt an LLM-based evaluation approach. However, LLM evaluation has limitations when it comes to assessing the practical usefulness of review comments in real-world scenarios. Therefore, we also introduce a human evaluation approach.
@@ III-D1 LLM Evaluation
We follow the LLM-based DeepCRCEval method by Lu et al. [30], with some modifications: (1) merging some metrics to improve clarity; (2) instructing the LLM to rate each comment on a 1–5 scale according to five metrics (higher scores indicating better quality), instead of the original 10-point scale, since prior study [21] shows that a 5-point scale is easier to understand and apply accurately, avoids the coarseness of a 3-point scale, and provides greater stability and consistency than a 10-point scale; (3) adding simple scoring guidelines for each metric. For example, for the “Readability” metric, we provided the LLM with the following explanation: “In the 1-5 scoring system, a score of 5 means the text is completely clear and readable; a score of 3 means the expressed meaning can be basically understood; and a score of 1 means the text is nearly unreadable or incomprehensible.” We believe these scoring guidelines help calibrate scores across different evaluation samples and positively impact the consistency of the LLM. Table II shows the final metrics.
@@ III-D2 Human Evaluation
Sometimes, LLM-based evaluation may not effectively assess the usefulness of the generated review comments in real-world scenarios. Therefore, we also conduct a human evaluation. Following Mahajan et al. [33], we adjust their method for LLM-generated results (Table III): we remove the “Unavailable” category, redefine the other three to suit our goals, and add an “Uncertain” category. This is because, in our evaluation, all test data have corresponding generated review comments, so the “Unavailable” case does not apply. However, given the limited knowledge of the repository and PR context, these review comments could be difficult to evaluate, leading to “Uncertain” cases. Specifically, since the data comes from 1,807 open-source projects, we cannot fully understand every project’s code requirements, organizational structure, or other contextual details. Therefore, even if we determine that the LLM-generated code review comments are generally useful, without factual grounding, we cannot be certain whether they would be accepted within the specific project context of the reviewed code. We categorize each generated review comment as “Instrumental”, “Helpful”, “Uncertain”, or “Misleading”, and use three metrics to evaluate quality.
@@ IV-A RQ1: Effectiveness of LAURA
In addition, in our experiments, the performance of directly using ChatGPT-4o for code review generation even surpasses that of CodeReviewer, which contrasts with the results of Tufano et al. [52]. We attribute this to advances in LLMs since ChatGPT-3, dataset differences, and varied evaluation methods. Our dataset, possibly containing more long code diffs, may have overwhelmed pretrained models. Additionally, Tufano et al. evaluated the three traditional methods as a whole, which also leads to differences in results.
@@ VI-A Code Review Automation
Early studies mainly focused on reviewer recommendation, using methods such as file path analysis [49], traceability graphs [47], load balancing [3], collaborative filtering [8], and balancing expertise and workload [35]. Other approaches considered cross-project collaboration [23] and composite analyses of file paths, project details, and author info [38]. Some research also aimed to predict which parts of the submitted code would need comments [19].
Code review automation has advanced to include tasks like review necessity prediction, comment generation, and code refinement. Notable works include T5CR [53], CodeReviewer [27], and CommentFinder [20]. T5CR, based on the Text-To-Text Transfer Transformer (T5) model [40], handles method-level code for pre-review optimization, comment generation, and post-review refinement. CodeReviewer uses T5 architecture with CodeT5 initialization [58], processes line-level diffs, and supports review necessity prediction, comment generation, and code refinement across nine languages. CommentFinder is an information retrieval (IR)-based method that uses Bag of Words (BoW) and similarity measures to find similar code snippets and recommend relevant review comments.
So far, several works have attempted to apply LLMs to code review automation tasks. Lu et al. [32] used the LLaMA model for review necessity prediction, comment generation, and code refinement, achieving results comparable to pre-trained models via parameter-efficient fine-tuning. Guo et al. [18] showed that ChatGPT-3.5 outperformed CodeReviewer in post-review refinement. Pornprasit et al. [39] explored ChatGPT-3.5’s performance in both pre- and post-review refinement, analyzing few-shot learning effects. Tufano et al. [52] compared ChatGPT-3 with three existing methods and found it achieved state-of-the-art performance in code refinement, but lagged in review generation. However, these studies provide only a limited exploration of LLMs and have not fully unlocked their potential for code review automation tasks.
@@ VI-B Large Language Model
Language models have evolved from early statistical models to neural, pre-trained models, and now to large language models (LLMs) with over 10B parameters, trained on vast datasets [65]. LLMs show superior performance and emergent capabilities [59], often matching or surpassing task-specific pre-trained models. Notable LLMs include ChatGPT-4o, which excels across benchmarks [37], and DeepSeek v3 [28] among open-source models. Our research builds on ChatGPT-4o and DeepSeek v3, applying the LAURA framework to both.
So far, LLMs have been widely used in software engineering tasks like vulnerability prediction [48, 67, 44] and commit message generation [12, 62, 26]. Many studies have explored ways to improve the performance of LLMs on specific tasks. Generally, LLMs can be further enhanced through methods such as augmenting input information [31], retrieval-augmented generation [14], and prompt engineering [61]. Augmenting input information is based on the intuitive idea that providing the model with richer contextual information helps it better understand the task and produce better results. Retrieval-augmented generation [25] enhances the model’s task awareness by incorporating knowledge from external databases, thereby strengthening the model’s generation capabilities. Prompt engineering [29] involves designing prompts for the model, guiding it through instructions on the actions it must perform with the given input, and ensuring that the LLMs achieve high performance in downstream tasks.